# DOPL: Direct Online Preference Learning for Restless Bandits with Preference Feedback

### v1

@@ 1 Introduction
In this paper, we consider the online Pref-RMAB, and show how to directly leverage the preference feedback to design a provably efficient RL algorithm for Pref-RMAB without explicit reward modeling. Specifically, we develop Direct Online Preference Learning (DOPL), an episodic RL algorithm that optimizes the same objective as the standard RMAB problem (cumulative reward maximization under an instantaneous constraint) but can efficiently explore the unknown environment, adaptively collect preference data in an online manner, and directly leverage the preference feedback to make decisions for Pref-RMAB.
@@
Third, to handle the instantaneous constraint in Pref-RMAB, we develop a low-complexity index policy (Section 3.3) for the DM to make decisions. Specifically, we first solve a relaxed problem, which turns out to be a linear programming (LP) in terms of occupancy measures, rather than Pref-RMAB itself. This is due to the fact that Pref-RMAB is computationally intractable like the standard RMAB (Papadimitriou & Tsitsiklis, 1994). Our key insight here is that we can define the objective of the LP for Pref-RMAB in terms of preference feedback directly. DOPL can therefore construct an index policy on top of the solutions to the LP. This is another key difference compared to the solutions (Whittle, 1988; Larrañaga et al., 2014; Bagheri & Scaglione, 2015; Verloop, 2016; Zhang & Frazier, 2021) to the standard RMAB, which heavily reply on the scalar reward.
We prove that DOPL achieves an $\tilde{\mathcal{O}}(\sqrt{T\ln T})$ regret, where $T$ is the time horizon. We believe that ours is the first work to formally characterize the regret for Pref-RMAB. Importantly, despite the fact that the DM in Pref-RMAB can only observe preference feedback, which contains arguably less information than scalar rewards, Pref-RMAB still achieves a sublinear regret as the standard RMAB, thanks to the novel design of online preference learning and direct index policy in DOPL.
@@ 2.2 Restless Multi-Armed Bandits with Preference Feedback
Similar to the standard RMAB, the goal of the DM in Pref-RMAB is still to maximize the long-term average reward as in (1) by activating a subset of $\mathcal{N}^{t}\subset\mathcal{N}$ arms at each time slot $t$ with $B=|\mathcal{N}^{t}|$. For ease of presentation, we refer to (1) as the optimization problem that the DM needs to solve in Pref-RMAB in the rest of this paper. Unlike the standard RMAB, the DM only observes a preference feedback in the form of the Bernoulli variable $\alpha(s_{m}^{t},s_{n}^{t})\sim\text{Ber}(\mathbf{F}_{m}^{n}(\sigma_{s_{m}^{t}},\sigma_{s_{n}^{t}}))$ between arms $(m,n)\in\mathcal{N}^{t}\times\mathcal{N}^{t}$ according to the underlying preference matrix $\mathbf{F}\in(0,1)^{N|\mathcal{S}|\times N|\mathcal{S}|}$. The probability that arm $m$ in state $s_{m}^{t}$ is preferred over arm $n$ in state $s_{n}^{t}$ is $\text{Pr}(\alpha(s_{m}^{t},s_{n}^{t})=1)=\mathbf{F}((m-1)|\mathcal{S}|+\sigma_{s_{m}^{t}},(n-1)|\mathcal{S}|+\sigma_{s_{n}^{t}})$. We assume that the preference feedback $\alpha(s_{m}^{t},s_{n}^{t})$ is drawn according to the widely-used Bradley-Terry (BT) model (Bradley & Terry, 1952), satisfying
@@ 3 Direct Online Preference Learning for Pref-RMAB
Lemma 1.
The underlying scalar reward $r_{n}(s)$ associated with arm $n$ in state $s$ can be expressed as
@@
Lemma 1 indicates that the scalar reward of arm $n$ in state $s$ can be expressed in terms of (i) the scalar reward of the reference arm $\star$ in a reference state $s_{\star}$, and (ii) the preference feedback between arm $n$ and the reference arm $\star$, which is drawn from the underlying BT model as in (2). Note that the reference arm $\star$ and reference state $s_{\star}$ can be any arm in $\mathcal{N}$ and any state in $\mathcal{S}$, but they will remain fixed in the system once selected. We will provide insights on how to select them in practice in our experimental evaluations in Section 5. More importantly, as we will show in Lemma 5, the objective of Pref-RMAB can be directly expressed in terms of preference feedback and is regardless of the scalar reward $r_{\star}(s_{\star})$ associated with the reference arm and the reference state.
Lemma 2.
Define the “preference-reference” term as $Q_{n}(s):=\ln\frac{\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}{1-\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}$. Since $r_{n}(s)\in[0,1],\forall s,n$, $Q_{n}(s)$ is monotonically increasing in $\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})$ and bounded as $Q_{n}(s)\in[-1,+1].$
@@
Similar to the scalar reward $r_{n}(s)$ in standard RMAB, Lemmas 1 and 2 indicates that the preference of arm $n$ in state $s$ in Pref-RMAB is well represented by $Q_{n}(s)$. Intuitively, the higher the probability (i.e., $\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})$ as defined in (2)) arm $n$ in state $s$ is preferred over the reference arm $\star$ and reference state $s_{\star},$ the larger the value of $Q_{n}(s)$ will be.
@@ 3.2 Online Preference Learning
From Lemma 2, the preference-reference term $Q_{n}(s):=\ln\frac{\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}{1-\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}$ relies on the reference arm $\star$ and reference state $s_{\star}$. Hence, to estimate the preference of arm $n$ in state $s$ or the value of $Q_{n}(s)$, DOPL only needs to learn one specific column of the preference matrix $\mathbf{F}$, i.e., $\mathbf{F}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})$. This corresponds to the preference between the reference state $s_{\star}$ of reference arm $\star$ with any arm $n\in\mathcal{N}$ in any state $s\in\mathcal{S}$. Since the pairwise preference feedback is a Bernoulli random variable, and the probability of the random variable being 1 is stored in preference matrix $\mathbf{F}$, an intuitive way is to estimate $\mathbf{F}$ with empirical dueling samples. Our online preference learning consists of four steps detailed as follows and the pseudocode is summarized in Algorithm 3 in Appendix C.3.
@@ 3.3 Direct Index Policy
It is known that solving the Pref-RMAB in (1) is computationally intractable even in the offline setting with scalar reward feedback (Papadimitriou & Tsitsiklis, 1994). To tackle this challenge and inspired by Whittle (Whittle, 1988), we first relax the instantaneous constraint, i.e., the constraint is satisfied on average rather than in each decision epoch. It turns out (Verloop, 2016; Zhang & Frazier, 2021; Xiong et al., 2022b) that this relaxed problem can be equivalently transformed into a LP (6)-(9) in terms of occupancy measures (Altman, 1999), where the occupancy measure $\mu^{\pi}$ of a stationary policy $\pi$ for $N$ controlled infinite-horizon MDPs is defined as the expected number of visits to each state-action pair $(s,a)$, i.e., $\mu^{\pi}=\{\mu_{n}(s,a)\triangleq\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi}\sum_{t=1}^{T}\mathds{1}(s_{n}^{t}=s,a_{n}^{t}=a):\forall n,s,a\}.$ It is obvious that the occupancy measure satisfies $\sum_{(s,a)}\mu_{n}(s,a)=1$, and hence $\mu_{n},\forall n\in\mathcal{N}$ is a probability measure. Unfortunately, we cannot solve this LP since we cannot directly observe the scalar rewards $r_{n}(s),\forall n,s$ in Pref-RMAB. To address this challenge, we first show that we can equivalently transform the LP (6)-(9) into a new LP in terms of preference feedback for Pref-RMAB.
@@
where $\omega^{k,*}=\{\omega_{n}^{k,*}(s,a,s^{\prime}),\forall n\in\mathcal{N}\}$. We present more details about this ELP in Appendix C.1. However, the optimal solution $\omega^{k,*}$ to (12) is not always feasible for Pref-RMAB due to the fact that the instantaneous constraint in Pref-RMAB must be satisfied in each decision epoch rather than on the average sense as in this ELP. Inspired by verloop2016asymptotically; zhang2021restless; xiong2022Nips, we further construct an index policy on top of $\omega^{k,*}$ that is feasible for Pref-RMAB. Specifically, the index assigned to arm $n$ in state $s_{n}^{t}=s$ at time $t$ is defined as $\mathcal{I}^{k}_{n}(s):=\frac{\sum_{s^{\prime}\in\mathcal{S}}\omega_{n}^{k,*}(s,1,s^{\prime})}{\sum_{a\in\mathcal{A},s^{\prime}\in\mathcal{S}}\omega_{n}^{k,*}(s,a,s^{\prime})}$. We call this the direct index since it is constructed by solving a LP in preference feedback directly. We then rank all arms according to their direct indices in a non-increasing order, and activate the set of $B$ highest indexed arms at each decision epoch. We denote the resultant direct index policy as $\pi^{k}$ and execute it in episode $k.$
@@ 4 Theoretical Guarantee
$Reg(T)$ $\leq\tilde{\mathcal{O}}\left(\left(c_{1}+\frac{2B}{1-D}\right)\sqrt{2NT|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\right),$ (13)
@@
The regret contains two terms. The first term with coefficient $c_{1}$ is the regret due to the online preference learning (Section 3.2). The second term with coefficient $2B/(1-D)$ is caused by executing the direct index policy for Pref-RMAB (to satisfy the instantaneous constraint) (Section 3.3). Clearly, the regret of DOPL is in the order of $\tilde{O}(\sqrt{T\ln T})$, which matches that of the standard RMAB with scalar rewards (Ortner et al., 2012; Wang et al., 2020; Xiong et al., 2022a). However, comparing to the standard RMAB with scalar rewards, our Pref-RMAB is a harder problem in which the DM can only observe preference feedback, which arguably contains less information. This challenging setting thus requires us to develop an effective online preference learning algorithm to handle preference feedback, as well as a new index policy that can directly leverage the preference feedback to make decisions.
@@
With probability $1-2\epsilon$, we have $Term_{1}\leq c_{1}\sqrt{2NT|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}.$
@@
With probability $1-2\epsilon$, we have $Term_{2}\leq\frac{2B}{1-D}\sqrt{2NT|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}.$
@@ 5 Experiments
In this section, we evaluate the efficacy of DOPL for Pref-RMAB via real-world applications. The experiment code is available at https://anonymous.4open.science/r/dopo-04C5.
@@ D.4 Proof of Lemma 4
To determine whether $f(x,y)$ is Lipschitz continuous over the interval $\left[\frac{1}{1+e},\frac{e}{1+e}\right]$, we need to analyze the partial derivatives of $f(x,y)$ and check if they are bounded within this interval. We start by calculating the partial derivates of $f$ with respect to $x$ and $y$:
@@ D.5 Proof of Lemma 5
According to Lemma 1, any reward $r_{n}(s)$ can be equivalently denoted as $r_{\star}(s_{\star})+Q_{n}(s)$. Then we can rewrite the LP in (6) as
@@ E.1 Proof of Lemma 6
It is straightforward to show that the LP in (6)-(9) is equivalent to the original RMAB problem in (1) if the "hard" activation constraint in (1) is relaxed to an averaged one as $\liminf\limits_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\sum_{n=1}^{N}A_{n}^{t}\right]\leq B$. To prove Lemme 11, it is sufficient to show that the relaxed problem achieves no less average reward than the original problem in (1). The proof is straightforward since the relaxed constraint expands the feasible region of the original problem in (1). Denote the feasible region of the original problem as

### v2

@@ 1 Introduction
In this paper, we consider the online Pref-RMAB, and show how to directly leverage the preference feedback to design a provably efficient RL algorithm for Pref-RMAB without explicit reward modeling. Specifically, we develop Direct Online Preference Learning (DOPL), an episodic RL algorithm that optimizes the same objective as the standard RMAB problem but can efficiently explore the unknown environment, adaptively collect preference data in an online manner, and directly leverage the preference feedback to make decisions for Pref-RMAB.
@@
Third, to handle the instantaneous constraint in Pref-RMAB, we develop a low-complexity index policy (Section 3.3) for the DM to make decisions. Specifically, we first solve a relaxed problem, which turns out to be a linear programming (LP) in terms of occupancy measures, rather than Pref-RMAB itself. Our key insight here is that we can define the objective of the LP for Pref-RMAB in terms of preference feedback directly. DOPL can therefore construct an index policy on top of the solutions to the LP. This is another key difference compared to the solutions (Whittle, 1988; Larrañaga et al., 2014; Bagheri & Scaglione, 2015; Verloop, 2016; Zhang & Frazier, 2021) to the standard RMAB, which heavily rely on the scalar reward.
We prove that DOPL achieves an $\tilde{\mathcal{O}}((2c_{1}+\frac{4B}{1-D})\sqrt{2NT|\mathcal{S}|\ln{4|\mathcal{S}||\mathcal{A}|NT}/{\epsilon}})$ regret, where $T$ is the time horizon, $\mathcal{S}$ is the state space, $D$ is the ergodicity coefficient and $c_{1},\epsilon$ are some constants (see details in Section 4). We believe that ours is the first work to formally characterize the regret for Pref-RMAB. Importantly, the two terms in this regret clearly maps to the aforementioned design of DOPL, i.e., the first term with coefficient $c_{1}$ is the regret due to the online preference learning and the second term with coefficient $4B/(1-D)$ is caused by executing the direct index policy for Pref-RMAB. Furthermore, despite the fact that the DM in Pref-RMAB can only observe preference feedback, which contains arguably less information than scalar rewards, Pref-RMAB still achieves a sublinear regret as the standard RMAB, thanks to the novel design of online preference learning and direct index policy in DOPL.
@@ 2.2 Restless Multi-Armed Bandits with Preference Feedback
Similar to the standard RMAB, the goal of the DM in Pref-RMAB is still to maximize the long-term average reward as in (1) by activating a subset of $\mathcal{N}^{t}\subset\mathcal{N}$ arms at each time slot $t$ with $B=|\mathcal{N}^{t}|$. For ease of presentation, we refer to (1) as the optimization problem that the DM needs to solve in Pref-RMAB in the rest of this paper. Unlike the standard RMAB, the DM only observes a preference feedback in the form of the Bernoulli variable $\alpha(s_{m}^{t},s_{n}^{t})\sim\text{Ber}(\mathbf{F}_{m}^{n}(\sigma_{s_{m}^{t}},\sigma_{s_{n}^{t}}))$ between arms $(m,n)\in\mathcal{N}^{t}\times\mathcal{N}^{t}$ according to the underlying preference matrix $\mathbf{F}\in(0,1)^{N|\mathcal{S}|\times N|\mathcal{S}|}$. The probability [^1] This probability measure can be extended to cases where $r_{n}(s,0)\neq 0$ by incorporating the action-dependence of operator $\mathbf{F}$. that arm $m$ in state $s_{m}^{t}$ is preferred over arm $n$ in state $s_{n}^{t}$ is $\text{Pr}(\alpha(s_{m}^{t},s_{n}^{t})=1)=\mathbf{F}((m-1)|\mathcal{S}|+\sigma_{s_{m}^{t}},(n-1)|\mathcal{S}|+\sigma_{s_{n}^{t}})$. We assume that the preference feedback $\alpha(s_{m}^{t},s_{n}^{t})$ is drawn according to the widely-used Bradley-Terry (BT) model (Bradley & Terry, 1952), satisfying
@@ 3 Direct Online Preference Learning for Pref-RMAB
Proposition 1.
The underlying scalar reward $r_{n}(s)$ associated with arm $n$ in state $s$ is expressed as
@@
Proposition 1 indicates that the scalar reward of arm $n$ in state $s$ can be expressed in terms of (i) the scalar reward of the reference arm $\star$ in a reference state $s_{\star}$, and (ii) the preference feedback between arm $n$ and the reference arm $\star$, which is drawn from the underlying BT model as in (2). For simplicity, define the “preference-reference” term as $Q_{n}(s):=\ln\frac{\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}{1-\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}$. Note that the reference arm $\star$ and reference state $s_{\star}$ can be any arm in $\mathcal{N}$ and any state in $\mathcal{S}$, but they will remain fixed in the system once selected. We will provide insights on how to select them in practice in our experimental evaluations in Section 5. More importantly, as we will show in Lemma 3, the objective of Pref-RMAB can be directly expressed in terms of preference feedback and is regardless of the scalar reward $r_{\star}(s_{\star})$ associated with the reference arm and the reference state.
Proposition 2.
Since $r_{n}(s)\in[0,1],\forall s,n$, $Q_{n}(s)$ is monotonically increasing in $\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})$ and bounded as $Q_{n}(s)\in[-1,+1].$
@@
Similar to the scalar reward $r_{n}(s)$ in standard RMAB, Propositions 1 and 2 indicate that the preference of arm $n$ in state $s$ in Pref-RMAB is well represented by $Q_{n}(s)$. Intuitively, the higher the probability (i.e., $\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})$ as defined in (2)) arm $n$ in state $s$ is preferred over the reference arm $\star$ and reference state $s_{\star},$ the larger the value of $Q_{n}(s)$ will be.
@@ 3.2 Online Preference Learning
From Prop 2, the preference-reference term $Q_{n}(s):=\ln\frac{\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}{1-\mathbf{F}_{n}^{\star}(\sigma_{s},\sigma_{s_{\star}})}$ relies on the reference arm $\star$ and reference state $s_{\star}$. Hence, to estimate the preference of arm $n$ in state $s$ or the value of $Q_{n}(s)$, DOPL only needs to learn one specific column of the preference matrix $\mathbf{F}$, i.e., $\mathbf{F}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})$. This corresponds to the preference between the reference state $s_{\star}$ of reference arm $\star$ with any arm $n\in\mathcal{N}$ in any state $s\in\mathcal{S}$. Since the pairwise preference feedback is a Bernoulli random variable, and the probability of the random variable being 1 is stored in preference matrix $\mathbf{F}$, an intuitive way is to estimate $\mathbf{F}$ with empirical dueling samples. Our online preference learning consists of four steps detailed as follows and the pseudocode is summarized in Algorithm 3 in Appendix C.3.
@@ 3.3 Direct Index Policy
It is known that solving the Pref-RMAB in (1) is computationally intractable even in the offline setting with scalar reward feedback (Papadimitriou & Tsitsiklis, 1994). To tackle this challenge and inspired by Whittle (Whittle, 1988), we first relax the instantaneous constraint, i.e., the constraint is satisfied on average rather than in each decision epoch. It turns out (Verloop, 2016; Zhang & Frazier, 2021; Xiong et al., 2022b) that this relaxed problem can be equivalently transformed into a LP (6)-(9) in terms of occupancy measures (Altman, 1999). Specifically, the occupancy measure $\mu_{n}(s,a)\triangleq\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi}\sum_{t=1}^{T}\mathds{1}(s_{n}^{t}=s,a_{n}^{t}=a)$ is the expected number of visits to each state-action pair $(s,a)$ of arm $n$, and $\mu^{\pi}$ of a stationary policy $\pi$ for $N$ controlled infinite-horizon MDPs is defined as $\mu^{\pi}=\{\mu_{n}(s,a):\forall n,s,a\}.$ It is obvious that the occupancy measure satisfies $\sum_{(s,a)}\mu_{n}(s,a)=1$, and hence $\mu_{n},\forall n\in\mathcal{N}$ is a probability measure. Unfortunately, we cannot solve this LP since we cannot directly observe the scalar rewards $r_{n}(s),\forall n,s$ in Pref-RMAB. To address this challenge, we first show that we can equivalently transform the LP (6)-(9) into a new LP in terms of preference feedback for Pref-RMAB.
@@
where $\omega^{k,*}=\{\omega_{n}^{k,*}(s,a,s^{\prime}),\forall n\in\mathcal{N}\}$. We present more details about this ELP in Appendix C.1. However, the optimal solution $\omega^{k,*}$ to (12) is not always feasible for Pref-RMAB due to the fact that the instantaneous constraint in Pref-RMAB must be satisfied in each decision epoch rather than on the average sense as in this ELP. Inspired by verloop2016asymptotically; zhang2021restless; xiong2022Nips, we further construct an index policy on top of $\omega^{k,*}$ that is feasible for Pref-RMAB. Specifically, the index assigned to arm $n$ in state $s_{n}^{t}=s$ at time $t$ is defined as $\mathcal{I}^{k}_{n}(s):=\frac{\sum_{s^{\prime}\in\mathcal{S}}\omega_{n}^{k,*}(s,1,s^{\prime})}{\sum_{a\in\mathcal{A},s^{\prime}\in\mathcal{S}}\omega_{n}^{k,*}(s,a,s^{\prime})}$. We call this the direct index since it is constructed by solving a LP in preference feedback directly. We then rank all arms according to their direct indices in a non-increasing order, and activate the set of $B$ highest indexed arms at each decision epoch. We denote the resultant direct index policy as $\pi^{k}$ and execute it in episode $k.$ Asymptotic optimality of this index policy relies on the existence of a uniform global attractor (Gast et al., 2024), a standard assumption in the RMAB literature.
@@ 4 Theoretical Guarantee
$Reg(T)$ $\leq\tilde{\mathcal{O}}\left(\left(2c_{1}+\frac{4B}{1-D}\right)\sqrt{2NT|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\right),$ (13)
@@
The regret contains two terms. The first term with coefficient $c_{1}$ is the regret due to the online preference learning (Section 3.2). The second term with coefficient $4B/(1-D)$ is caused by executing the direct index policy for Pref-RMAB (to satisfy the instantaneous constraint) (Section 3.3). Clearly, the regret of DOPL is in the order of $\tilde{O}(\sqrt{T\ln T})$, which matches that of the standard RMAB with scalar rewards (Ortner et al., 2012; Wang et al., 2020; Xiong et al., 2022a). However, comparing to the standard RMAB with scalar rewards, our Pref-RMAB is a harder problem in which the DM can only observe preference feedback, which arguably contains less information. This challenging setting thus requires us to develop an effective online preference learning algorithm to handle preference feedback, as well as a new index policy that can directly leverage the preference feedback to make decisions.
@@
With probability $1-2\epsilon$, we have $Term_{1}\leq 2c_{1}\sqrt{2N^{2}T|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}.$
@@
With probability $1-2\epsilon$, we have $Term_{2}\leq\frac{4B}{1-D}\sqrt{2N^{2}T|\mathcal{S}|\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}.$
@@ 5 Experiments
In this section, we evaluate the efficacy of DOPL for Pref-RMAB via real-world applications. The experiment code is available at https://github.com/flash-36/DOPL.
@@ 6 Conclusion
Acknowledgments
Research was funded in part by grants NSF CNS 2148309, 2312978 and 2337914. All all opinions expressed are those of the authors, and need not represent those of the sponsoring agencies
@@ D.4 Proof of Lemma 2
To determine whether $f(x,y)$ is Lipschitz continuous over the interval $\left[\frac{1}{1+e},\frac{e}{1+e}\right]$, we need to analyze the partial derivatives of $f(x,y)$ and check if they are bounded within this interval. Let us start by calculating the partial derivatives of $f$ with respect to $x$ and $y$:
@@ D.5 Proof of Lemma 3
According to Proposition 1, any reward $r_{n}(s)$ can be equivalently denoted as $r_{\star}(s_{\star})+Q_{n}(s)$. Then we can rewrite the LP in (6) as
@@ E.1 Proof of Lemma 4
It is straightforward to show that the LP in (6)-(9) is equivalent to the original RMAB problem in (1) if the "hard" activation constraint in (1) is relaxed to an averaged one as $\liminf\limits_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\sum_{n=1}^{N}A_{n}^{t}\right]\leq B$. To prove Lemme 9, it is sufficient to show that the relaxed problem achieves no less average reward than the original problem in (1). The proof is straightforward since the relaxed constraint expands the feasible region of the original problem in (1). Denote the feasible region of the original problem as